OpenStudy (anonymous):

Find the derivative, r'(t), of the vector function. r(t) = at cos 2ti + b sin^5 tj + c cos^5 tk

ganeshie8 (ganeshie8):

take the derivative of each component

ganeshie8 (ganeshie8):

\(\large r'(t) = (at \cos 2t)' i + (b \sin^5 t)' j + (c \cos^5 t)' k\)

OpenStudy (anonymous):

(acos(2t)-2atsin(2t))i+(5bcos(t)sin^2(t))j-(5csin(t)cos^2(t))k ??

ganeshie8 (ganeshie8):

i see few mistakes/typoes check again :)

OpenStudy (anonymous):

sin^4 ???:S

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

and cos ^4 right:S

ganeshie8 (ganeshie8):

yes thats it !

OpenStudy (anonymous):

thanks:)

ganeshie8 (ganeshie8):

np :)